OpenStudy (yisselluna):

Solve the following system of equations by the substitution method. x - y = 0 x - y - 2 = 0 What is the solution set? {(1, -1)} {(-1, 1)} ∅

OpenStudy (yisselluna):

i think its the first one? but not sure

OpenStudy (dallascowboys88):

yep

OpenStudy (ben00):

It can be anything. 1 - 1 = 0 -1 + 1 = 0

OpenStudy (dallascowboys88):

Yep that too but yes ur right

OpenStudy (ben00):

Maybe (1, -1)

OpenStudy (ben00):

Please give medal to correct answers.

OpenStudy (dallascowboys88):

exactly the first one like i said

OpenStudy (dallascowboys88):

wow LIKE I SAID THE FIRST ONE WTHECK DUDE I Said the correct answer before you both answers are right

OpenStudy (ben00):

Well then

OpenStudy (will.h):

IT IS NOT the 1st one

OpenStudy (will.h):

i was trying to explain but clearly you looking for direct answers which i won't give Bye

OpenStudy (yisselluna):

lmaooo okay byeeeeeeeee